# QoE Maximization for Multiple-UAV-Assisted Multi-Access Edge Computing: An Online Joint Optimization Approach

### v1

@@ 6 Simulation Results
6.1 Simulation Setup
We first describe the settings related to the simulation experiments, including the scenario, parameters, comparative approaches, and performance evaluation metrics.
6.1.1 Scenario
We consider a multiple-UAV-assisted MEC system, where one LUAV and four SUAVs are deployed to collaboratively provide edge computing services for $60$ ground UDs in a $1,000\times 1,000\ \text{m}^{2}$ rectangular service area. The system timeline is divided into $100$ equal time slots, and the duration of each time slot is set to $\Delta t=1\ \text{s}$.
6.1.2 Parameters
TABLE II: Simulation Parameters
@@ 6.1.2 Parameters
$\varpi$ Energy consumption per unit CPU cycle of SUAVs $8.2\times 10^{-27}$[34]
$C_{1}$, $C_{2}$, $C_{3}$, $C_{4}$ UAV propulsion power consumption parameters 80, 22, 263.4, 0.0092 [43]
$\bar{E}_{n}$ Energy budget per time slot for SUAV $n$ 220 J
@@
For the LUAV, the fixed horizontal position and altitude are set as $\mathbf{q}_{u}=[500,500]$ and $H_{u}=300\ \text{m}$, respectively. For the SUAVs, the fixed altitude is set as $H=100\ \text{m}$, and the initial positions are set as $\mathbf{q}_{1}^{\text{ini}}=[100,100]$, $\mathbf{q}_{2}^{\text{ini}}=[100,900]$, $\mathbf{q}_{3}^{\text{ini}}=[900,900]$, and $\mathbf{q}_{4}^{\text{ini}}=[900,100]$, respectively. For the UDs, the computing capacity of each UD is randomly taken from $\{1,1.5,2\}\ \text{GHz}$. The default values for the remaining parameters can be found in Table II.
6.1.3 Comparative Approaches
@@
  • Entire offloading (EO) [54]: All UDs offload their tasks to aerial servers for execution. This approach does not consider local computing.
@@
  • DDPG-based joint optimization (DDPG-JO) [32]: The decisions of task offloading, resource allocation, and trajectory planning are decided by the DDPG algorithm.
  • PPO-based joint optimization (PPO-JO) [33]: The decisions of task offloading, resource allocation, and trajectory planning are determined by the PPO algorithm.
6.1.4 Performance Metrics
@@ 6.2 Evaluation Results
In this section, we first evaluate the system performance of the proposed OJTRTA over time with default parameters. Subsequently, we compare the impacts of different parameters on the performance of the proposed OJTRTA and the benchmark approaches.
6.2.1 Online Offloading Performance Evaluation
@@
Figs. 2(a), 2(b), 2(c) and 2(d) show the dynamics of time-averaged UD cost, average task completion latency, cumulative UD energy consumption, and time-averaged SUAV energy consumption, respectively among the seven approaches. It can be observed that the time-averaged UD cost exhibits some fluctuations over time. This is primarily attributed to the time-varying computational demands of UDs. Additionally, OCQ slightly outperforms the proposed OJTRTA in terms of time-averaged UD cost, average task completion latency, and overall UD energy consumption. This is because OCQ does not consider the energy constraints of SUAVs and thus consumes more SUAV energy. Furthermore, the proposed OJTRTA outperforms EO, ERA, and FLP in terms of time-averaged UD cost and average task completion latency. This is due to the combined local and edge offloading strategy, optimal resource allocation strategy, and trajectory planning strategy of OJTRTA. Moreover, compared to DRL-based approaches, i.e., DDPG-JO and PPO-JO, the proposed OJTRTA exhibits significant advantages in terms of time-averaged UD cost, average task completion delay, and cumulative UD energy consumption. This further exemplifies the effectiveness of the proposed approach. Finally, as shown in Fig. 2(d), the proposed OJTRTA ensures long-term energy constraints satisfaction for SUAVs under the real-time guidance of the Lyapunov-based energy queue, which is consistent with the analysis in Theorem 10. In conclusion, the set of simulation results demonstrates the effectiveness of the proposed approach in achieving better overall performance while satisfying the long-term energy consumption constraint of SUAVs.
6.2.2 Impact of Parameters
Figure 2: System performance with respect to time slots. (a) Time-averaged UD cost. (b) Average task completion latency. (c) Cumulative UD energy consumption. (d) Time-averaged SUAV energy consumption.
Figure 3: System performance with respect to different numbers of UDs. (a) Time-averaged UD cost. (b) Average task completion latency. (c) Cumulative UD energy consumption. (d) Time-averaged SUAV energy consumption.
Impact of UD Numbers. Figs. 3(a), 3(b), 3(c), and 3(d) illustrate the impact of varying numbers of UDs on the time-averaged UD cost, average task completion delay, cumulative UD energy consumption, and time-averaged SUAV energy consumption, respectively. It can be observed that the time-averaged UD cost, cumulative UD energy consumption and time-averaged SUAV energy consumption of all approaches show an upward trend as the number of UDs increases, since more computation tasks need to be processed. Furthermore, as the number of UDs increases, EO exhibits the poorest performance in terms of average task completion latency, while demonstrating optimal performance in cumulative UD energy consumption. This is primarily due to EO offloading all tasks to edge servers, resulting in lower UD energy consumption but higher server loads. Additionally, compared to the proposed approach, FLP shows inferior performance in terms of average task completion latency and cumulative UD energy consumption. This highlights the importance of optimizing the trajectory of UAVs. Besides, when the number of UDs reaches 100, OCQ and PPO-JO slightly outperform the proposed approach in terms of time-averaged UD cost. This can be attributed to the fact that OCQ and PPO-JO do not consider SUAV energy consumption constraints. As shown in Fig. 3(d), OCQ and PPO-JO exhibit higher time-averaged SUAV energy consumption.
Lastly, the proposed OJTRTA approach demonstrates superior performance in terms of time-averaged UD cost and average task completion latency compared to other approaches, while also exhibiting favorable performance in terms of cumulative UD energy consumption and time-averaged SUAV energy consumption. Specifically, compared with EO, ERA, FLP, and DDPG-JO, the proposed approach can respectively reduce the time-averaged UD cost by 54.7%, 2.3%, 1.7%, 8.8% and can respectively reduce the average task completion latency 58.1%, 0.4%, 2.6%, 4.2% in the relative dense scenario ($M=100$). In conclusion, the set of simulation results indicates that the proposed OJTRTA has better scalability with an increasing number of UDs.
@@ 6.2.2 Impact of Parameters
Figure 4: System performance with respect to different task data sizes. (a) Time-averaged UD cost. (b) Average task completion latency. (c) Cumulative UD energy consumption. (d) Time-averaged SUAV energy consumption.
@@
Figure 5: System performance with respect to different SUAV computing resources. (a) Time-averaged UD cost. (b) Average task completion latency. (c) Cumulative UD energy consumption. (d) Time-averaged SUAV energy consumption.
@@
Impact of Task Data Size. Figs. 4(a), 4(b), 4(c), and 4(d) show the impact of task data sizes on the time-averaged UD cost, average task completion latency, cumulative UD energy consumption, and time-averaged SUAV energy consumption. It can be seen that with the increase in task data sizes, there is an upward trend in terms of the time-averaged UD cost, average task completion latency, cumulative UD energy consumption, and time-averaged SUAV energy consumption. This is expected as the larger task data size leads to higher overheads on computing, communication, and energy consumption for UDs and UAVs. In addition, EO exhibits a significant growth trend in terms of the time-averaged UD cost and average task completion latency compared to other approaches. This is due to the increased competition for the limited computational and communication resources of aerial servers caused by the entire offloading of EO as the task data size increases.
@@
Finally, compared with EO, ERA, FLP, DDPG-JO and PPO-JO, the proposed OJTRTA achieves performance improvements of approximately 27.6%, 9.1%, 3.4%, 9.3%, and 6.8% in terms of time-averaged UD cost, as well as 30%, 9.2%, 2.8%, 7.5%, and 3.3% in terms of average task completion latency when the task data size reaches 1 Mb. The set of simulation results indicates that the proposed OJTRTA is able to adapt to the heavy-loaded scenarios with overall superior performances.
@@
Impact of Computing Resources of SUAVs. Figs. 5(a), 5(b), 5(c), and 5(d) depict the impact of different SUAV computing resources on the time-averaged UD cost, average task completion delay, cumulative UD energy consumption, and time-averaged SUAV energy consumption. It can be observed that with the increase of SUAV computing resources, all approaches show a decreasing trend in terms of the time-averaged UD cost and average task completion latency, and EO and OJTRTA demonstrate gradually decreasing performance improvements. The reasons can be explained as follows. The increase of SUAV computational resources provides more computing resource allocation for task execution, reducing the task execution latency. However, as SUAV computing resources further increase, communication resources and the energy constraints of SUAV become bottlenecks that limit the improvement of system performance. Furthermore, EO maintains nearly constant performance in terms of the cumulative UD energy consumption regardless of the variations in the computing resources of SUAVs. This is mainly because the entire offloading of EO primarily incurs transmission energy consumption for UDs, which is independent of the computing resources of SUAVs.
@@
6.2.3 UAV Trajectory Planning
Figure 6: Trajectories of the UDs and SUAVs.
@@ 6.2.3 UAV Trajectory Planning
7 Conclusion
In this work, we studied the task offloading, resource allocation, and UAV trajectory planning in an energy-constrained UAV-enabled MEC system. A JTRTOP was formulated to maximize the QoE of all UDs while satisfying the UAV energy consumption constraint. Since the JTRTOP is future-dependent and NP-hard, we proposed the OJTRTA to solve the problem. Specifically, the future-dependent JTRTOP was first transformed into the PROP by using Lyapunov optimization methods. Furthermore, a two-stage optimization algorithm was proposed to solve the PROP. Simulation results show that OJTRTA outperforms the comparative approaches in terms of time-averaged UD cost while meeting the SUAV energy consumption constraint. Furthermore, JTRTOP exhibits superior adaptability in heavy-loaded scenarios and demonstrates good scalability with an increasing number of MDs. In the future, we will integrate satellite networks into our work to explore the integration of space-air-ground MEC networks.

### v2

@@ VI Simulation Results
VI-A Simulation Setup
TABLE I: Simulation Parameters
@@ VI-A Simulation Setup
$\varpi$ Energy consumption per unit CPU cycle of SUAVs $8.2\times 10^{-9}$[42]
@@
VI-A1 Scenario
We consider a multiple-UAV-assisted MEC system, where one LUAV and four SUAVs are deployed to collaboratively provide edge computing services for $60$ ground UDs in a $1,000\times 1,000\ \text{m}^{2}$ rectangular service area. The system timeline is segmented into $100$ equal time slots, and the duration of each time slot is set to $\Delta t=1\ \text{s}$.
VI-A2 Parameters
For the LUAV, the fixed horizontal position and altitude are set as $\mathbf{q}_{u}=[500,500]$ and $H_{u}=300\ \text{m}$, respectively. For the SUAVs, the fixed altitude is set as $H=100\ \text{m}$, and the initial positions are set as $\mathbf{q}_{1}^{\text{ini}}=[100,100]$, $\mathbf{q}_{2}^{\text{ini}}=[100,900]$, $\mathbf{q}_{3}^{\text{ini}}=[900,900]$, and $\mathbf{q}_{4}^{\text{ini}}=[900,100]$, respectively. For the UDs, the computing capacity of each UD is randomly taken from $\{1,1.5,2\}\ \text{GHz}$. The default values for the remaining parameters can be found in Table I.
VI-A3 Comparative Approaches
@@
  • Entire offloading (EO): All UDs offload their tasks to aerial servers for execution. This approach does not consider local computing.
@@
  • DDPG-based joint optimization (DDPG-JO) [32]: The decisions of resource allocation, task offloading, and trajectory control are decided by the DDPG algorithm.
  • PPO-based joint optimization (PPO-JO) [7]: The decisions of resource allocation, task offloading, and trajectory control are determined by the PPO algorithm.
VI-A4 Performance Metrics
Figure 3: The impact of time slots on system performance. (a) Time-averaged UD cost. (b) Average task completion latency. (c) Cumulative UD energy consumption. (d) Time-averaged SUAV energy consumption.
Figure 4: The impact of the numbers of UDs on system performance. (a) Time-averaged UD cost. (b) Average task completion latency. (c) Cumulative UD energy consumption. (d) Time-averaged SUAV energy consumption.
@@ VI-B Evaluation Results
In this section, we first evaluate the performance of the proposed OJTRTA over time with default parameters. Then, we compare the impacts of different parameters on the performance of OJTRTA in comparison with benchmark approaches.
VI-B1 Online Offloading Performance Evaluation
@@
Figs. 3(a), 3(b), 3(c), and 3(d) show the dynamics of time-averaged UD cost, average task completion latency, cumulative UD energy consumption, and time-averaged SUAV energy consumption, respectively among the seven approaches. It can be observed that the time-averaged UD cost exhibits some fluctuations over time. This is primarily attributed to the time-varying computational demands of UDs. Additionally, OCQ slightly outperforms the proposed OJTRTA in terms of time-averaged UD cost, average task completion latency, and overall UD energy consumption. This is because OCQ does not consider the energy constraints of SUAVs and thus consumes more SUAV energy. Furthermore, the proposed OJTRTA outperforms EO, ERA, and FLP in terms of time-averaged UD cost and average task completion latency. This is due to the combined local and edge offloading strategy, optimal resource allocation strategy, and trajectory control strategy of OJTRTA. Moreover, compared to DRL-based approaches, i.e., DDPG-JO and PPO-JO, the proposed OJTRTA exhibits significant advantages in terms of time-averaged UD cost, average task completion delay, and cumulative UD energy consumption. This further exemplifies the effectiveness of the proposed approach. Finally, as shown in Fig. 3(d), the proposed OJTRTA ensures that the long-term energy constraints of SUAVs are satisfied, which is consistent with the analysis in Theorem 12. In conclusion, the set of simulation results demonstrates the effectiveness of the proposed approach in achieving better overall performance while satisfying the long-term energy consumption constraints of SUAVs.
@@
Figure 5: The impact of task data sizes on system performance. (a) Time-averaged UD cost. (b) Average task completion latency. (c) Cumulative UD energy consumption. (d) Time-averaged SUAV energy consumption.
@@
Figure 6: The impact of SUAV computing resources on system performance. (a) Time-averaged UD cost. (b) Average task completion latency. (c) Cumulative UD energy consumption. (d) Time-averaged SUAV energy consumption.
VI-B2 Impact of UD Numbers
Figs. 4(a), 4(b), 4(c), and 4(d) illustrate the impact of varying numbers of UDs on the time-averaged UD cost, average task completion delay, cumulative UD energy consumption, and time-averaged SUAV energy consumption, respectively. It can be observed that the time-averaged UD cost, cumulative UD energy consumption and time-averaged SUAV energy consumption of all approaches show an upward trend as the number of UDs increases, since more computation tasks need to be processed. Furthermore, as the number of UDs increases, EO exhibits the poorest performance in terms of average task completion latency, while demonstrating optimal performance in cumulative UD energy consumption. This is primarily due to EO offloading all tasks to edge servers, resulting in lower UD energy consumption but higher server loads. Additionally, compared to the proposed approach, FLP shows inferior performance in terms of average task completion latency and cumulative UD energy consumption. This highlights the importance of optimizing the trajectory of UAVs. Besides, when the number of UDs reaches 100, OCQ and PPO-JO slightly outperform the proposed approach in terms of time-averaged UD cost. This can be attributed to the fact that OCQ and PPO-JO do not consider SUAV energy consumption constraints. As shown in Fig. 4(d), OCQ and PPO-JO exhibit higher time-averaged SUAV energy consumption.
Lastly, the proposed OJTRTA approach demonstrates superior performance in terms of time-averaged UD cost and average task completion latency compared to other approaches, while also exhibiting favorable performance in terms of cumulative UD energy consumption and time-averaged SUAV energy consumption. Specifically, compared with EO, ERA, FLP, DDPG-JO, and PPO-JO, the proposed approach can respectively reduce the time-averaged UD cost by 55.6%, 3.2%, 2.1%, 9.4%, 14.6% and can respectively reduce the average task completion latency 58.7%, 1.2%, 3.1%, 4.5%, 15.4% in the relative dense scenario ($M=100$). In conclusion, the set of simulation results indicates that the proposed OJTRTA has better scalability with an increasing number of UDs.
VI-B3 Impact of Task Data Size
@@
Figs. 5(a), 5(b), 5(c), and 5(d) show the impact of task data sizes on the time-averaged UD cost, average task completion latency, cumulative UD energy consumption, and time-averaged SUAV energy consumption. It can be seen that with the increase in task data sizes, there is an upward trend in terms of the time-averaged UD cost, average task completion latency, cumulative UD energy consumption, and time-averaged SUAV energy consumption. This is expected as the larger task data size leads to higher overheads on computing, communication, and energy consumption for UDs and UAVs. Moreover, EO exhibits a significant growth trend in terms of the time-averaged UD cost and average task completion latency compared to other approaches. This is due to the increased competition for the limited computational and communication resources of aerial servers caused by the entire offloading of EO as the task data size increases.
@@
Finally, compared with EO, ERA, FLP, DDPG-JO and PPO-JO, OJTRTA achieves performance improvements of approximately 28.4%, 9.7%, 3.8%, 9.8%, and 7.2% in terms of time-averaged UD cost, as well as 30.4%, 9.3%, 3.2%, 8.5%, and 3.9% in terms of average task completion latency when the task data size reaches 1 Mb. The set of simulation results indicates that the proposed OJTRTA is able to adapt to the heavy-loaded scenarios with overall superior performances.
VI-B4 Impact of Computing Resources of SUAVs
@@
Figs. 6(a), 6(b), 6(c), and 6(d) depict the impact of different SUAV computing resources on the time-averaged UD cost, average task completion delay, cumulative UD energy consumption, and time-averaged SUAV energy consumption. It can be observed that with the increase of SUAV computing resources, all approaches show a decreasing trend in terms of the time-averaged UD cost and average task completion latency, and EO and OJTRTA demonstrate gradually decreasing performance improvements. The reasons can be explained as follows. The increase of SUAV computational resources provides more computing resource allocation for task execution, reducing the task execution latency. However, as SUAV computing resources further increase, communication resources and the energy constraints of SUAV become bottlenecks that limit the improvement of system performance. Furthermore, EO maintains nearly constant performance in terms of the cumulative UD energy consumption regardless of the variations in the computing resources of SUAVs. This is mainly because the entire offloading of EO primarily incurs transmission energy consumption for UDs, which is independent of the computing resources of SUAVs.
@@
VI-B5 UAV Trajectory Control
@@
Figure 7: Trajectories of UDs and SUAVs.
VII Conclusion
In this work, we investigated the task offloading, resource allocation, and UAV trajectory control within a multi-UAV-assisted MEC system. A JTRTOP was formulated to maximize the QoE of all UDs while satisfying the energy and resource constraints of UAVs. Since the JTRTOP is future-dependent and NP-hard, we proposed the OJTRTA to solve the problem. Specifically, the future-dependent JTRTOP was first transformed into the PROP by using Lyapunov optimization methods. Furthermore, a two-stage optimization method was proposed to solve the PROP. Simulation results show that OJTRTA outperforms the comparative approaches in terms of time-averaged UD cost while meeting the SUAV energy consumption constraint. Furthermore, JTRTOP exhibits superior adaptability in heavy-loaded scenarios and demonstrates good scalability as the number of MDs increases. In the future, our efforts will focus on extending current research to space-air-ground integrated MEC networks. Specifically, the proposed multi-UAV-assisted MEC system heavily relies on the limited computational resources and energy supply of UAVs, which remains a critical bottleneck restricting the improvement of system performance. With the rapid deployment of low Earth orbit satellites, space-air-ground integrated networks demonstrate significant potential to substantially expand available network resources and the service coverage of MEC by aggregating the capabilities of heterogeneous network.